OpenStudy (anonymous):

Happy Feet Rentals charges $14 per day for snorkel gear and $16 per day to rent a wetsuit. Sam and Monica paid a total of $300 to rent a set of snorkel gear and a wetsuit each for a few days. Monica rented them for 2 days less than Sam. The following equation represents this situation, where x represents the number of days Sam rented a surfboard and a wetsuit. 14x + 16x + 14(x − 2) + 16(x − 2) = 300 What is the total amount Sam paid to rent the snorkel gear and wetsuit? $80 $150 $180 $240

OpenStudy (anonymous):

Distribute. 14x + 16x + 14(x − 2) + 16(x − 2) = 300 14x + 16x + 14x − 28 + 16x − 32 = 300 Combine like terms. 14x + 16x + 14x − 28 + 16x − 32 = 300 60x − 28 − 32 = 300 60x - 60 = 300 Solve for x. 60x - 60 = 300 60x = 360 x = 360 /60 x = ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@jazy

OpenStudy (anonymous):

Right! (:

OpenStudy (anonymous):

so now what

OpenStudy (anonymous):

cause thats not the answer

OpenStudy (anonymous):

Check out @chemwile exxplanation here: http://openstudy.com/study#/updates/4f52cddae4b019d0ebb0615c

OpenStudy (anonymous):

180 :)

OpenStudy (anonymous):

i thought so :)

OpenStudy (anonymous):

Yep, sorry I didn't finish explaining it myself (:

OpenStudy (anonymous):

its ok